ratio of the ages of a father and his son 10 years hence will be 5:3, while 10 years ago, it was 3:1. the ratio of the age of son to that of father today is
Answer by richwmiller(17219)

(f+10)/(s+10)=5/3,
(f-10)/(s-10)=3/1,
f = 40, s = 20, ratio today is 2/1